OpenStudy (anonymous):

What is the solution set of |2x - 3| - 10 = -1?

OpenStudy (anonymous):

The two little lines next the equation (for example): \[\left| X \right|\] literally mean the absolute value of x. The sign does not matter, x will always be positive. To solve the equation, you will have to do a (plus or minus) on the other side, like this: \[\left| 2x-3 \right|-10=-1\] \[\left| 2x-3 \right|=9\]\[2x-3= \pm9\] \[2x=3\pm9\] \[x= (3+9)\div2\] \[x= (3-9)\div2\] Solve for x, and you will have to different sets for x. I can prove it to you graphically too if you want :)